The distance the rocker far is 16 feet the first second 48 the second second 80 feet the third second what is the total distance the object will fall in six seconds
1
Expert's answer
2020-09-17T09:28:49-0400
"a_n=a_1+d(n-1)"

"d=48-16=32, a_1=16"

"a_6=16+32(6-1)=176\ ft"